OpenStudy (anonymous):

Use properties of logarithms and solve the equation. Check for extraneous solutions. log(base3)(x+2) + log(base3)(x-2)=2

OpenStudy (anonymous):

start with \[\log_3((x+2)(x-2))=2\] then rewrite in equivalent exponential form as \[(x+2)(x-2)=3^2\] and finally solve the quadratic

OpenStudy (anonymous):

you get \[x^2-4=9\] \[x^2=13\] and so \[x=\sqrt{13}\]

OpenStudy (anonymous):

Thank you!! So exponential form is the base of the first log, 3, to the power of what the initial equation equaled, 2?

OpenStudy (anonymous):

yes, after you use the property that \(\log(A)+\log(B)=\log(AB)\) to write as a single log

OpenStudy (anonymous):

for example \(\log_5(x+6)=4\iff x+6=5^4\)

OpenStudy (anonymous):

Awesome! Thanks for your help!!

OpenStudy (anonymous):

yw